b. Dent invests $74,000 View transaction list Journal entry worksheet Record the Dents admission and bonus. Note: Enter debits before credits. Date General Journal Debit Credit May 01, 2017 Record entry Clear entry View general journal
Answer #1

Ans-a- Total Capital of existing partners= $85,000+$70,000+$148,000= $303,000

Capital invested by Dent= $101,000

Total capital after new capital introduced by Dent= $303,000+$101,000= $404,000

Dent share in partner ship= 25%

Required share of capital by Dent= $404,000*25%= 101,000

Therefore, there is no adjustment on admission of Dent.

Journal Entries- Admission of Partner

Date General Journal Debit ($) Credit ($)
May 01, 2017 Cash A/c Dr. 101,000
Dent's Capital A/c 101,000
(To record the capital on admission of Dent in partnership)

Ans-b- Total Capital of existing partners= $85,000+$70,000+$148,000= $303,000

Capital invested by Dent= $74,000

Total capital after new capital introduced by Dent= $303,000+$74,000= $377,000

Dent share in partner ship= 25%

Required share of capital by Dent= $377,000*25%=$94,250

As capital invested by Dent is lower than required capital, therefore existing partner will give bonus to Dent to cover shortfall of $ 20,250 in proportion of 3:2:5.

Journal Entries- Admission of Partner

Date General Journal Debit ($) Credit ($)
May 01, 2017 Cash A/ c Dr. 74,000
Dent's Capital A/c 74,000
(To record the capital on admission of Dent in partnership)
May 01, 2017

Zeller's Capital A/c Dr.

($20,250*3/10)

6,075

Acker's Capital A/c Dr.

($20,250*2/10)

4,050

Benton's Capital A/c Dr.

($20,250*5/10)

10,125
Dent's Capital A/c 20,250
(To record the bonus distributed to new partner by existing partner)

Ans-c- Total Capital of existing partners= $85,000+$70,000+$148,000= $303,000

Capital invested by Dent= $132,000

Total capital after new capital introduced by Dent= $303,000+$132,000= $435,000

Dent share in partner ship= 25%

Required share of capital by Dent= $435,000*25%= $108,750

As capital invested by Dent is higher than required capital, bonus capital of $23,250($132,000-$108,750) introduced by Dent will be distributed in existing partner in the ratio of 3:2:5

Journal Entries- Admission of partner

Date General Journal Debit ($) Credit ($)
May 01, 2017 Cash A/c Dr. 132,000
Dent's Capital A/c 132,000
(To record the capital on admission of Dent in partnership)
May 01, 2017 Dent's Capital A/c Dr. 23,250

Zeller's Capital A/c

($23,250*3/10)

6,975

Acker's Capital A/c

($23,250*2/10)

4,650

Benton's Capital A/c

($23,250*5/10)

11,625
(To record the bonus distributed to existing partner by new partner)

Kindly give thumbsup if u like my answer...Thanks!!!